five defective electric toothbrushes were accedentally shipped to a drugstore by clean brush products along with 25 non-defective ones.
What is the pprobability that the first two toothbrushes are sold will be defective?
What is the probability the first two toothbrushes will not be defective?

Answer by Boreal(15235)

probability of first is (5/30) and second given the first defective is (4/29). Product is 20/870=2/87.
-
(25/30)(24/29)=600/870=20/29